OpenStudy (anonymous):

There are six chairs in a row. Person A picks one of the chairs at random. Then Person B picks a chair at random from the five chairs that remain. Find the chance that they pick chairs that are next to each other.

OpenStudy (anonymous):

The number of ways two people can pick tow chirs out of 6 is: \[V^6_2=6·5=30\] If chair 1 or 6 are taken there is only one chair next to them that are chair 2 and 5. That means to cases of adjacent chairs. For chairs 2,3,4 and 5, there are two chairs next to each one of them, that means 4·2 cases of adjacent chairs. Total cases of adjacent chairs is then 2+8=10. The probability is? That is for yourself

OpenStudy (gorv):

@sonmodi ans is 1/3